OpenStudy (anonymous):

Identify the 19th term of a geometric sequence where a1 = 14 and a9 = 358.80. Round the common ratio and 19th term to the nearest hundredth. I know that I need to find r first which I got r^8=25.63 then you have to do the square root of each side but I can't remember how to solve when you have the square root of a number and then a number on the outside of the square root in the little check mark part.

OpenStudy (amistre64):

if we divide it all by 14; we normalize it i believe

OpenStudy (amistre64):

a1 = 14 and a9 = 358.80 ------------------------ ; a1 = 1 and a9 = 25.6286 14 a9 = r^8 r = 25.6286^(1/8) = about 1.499997

OpenStudy (amistre64):

^8 and 8rt cancel each other

OpenStudy (amistre64):

8rt(n) = n^(1/8)

OpenStudy (anonymous):

I don't think thats right. I just need to know how to get an answer when I come across something like r^8=25.63 and then I have to sq each side I just don't know how to do it when there is a number on the outside in this case is 8 in the check mark part of the sq root symbol.

OpenStudy (anonymous):

\[\LARGE \begin{array}{l}{a_9} = {a_1} \cdot {r^{9 - 1}}\\358.80 = 14 \cdot {q^8}\\25.62 = {q^8}\\q = \pm 1.49\\\\{a_9} = {a_1} \cdot 1.49\\358.80 = 14 \cdot {(1.49)^8}\\358.80 = 14 \cdot 24.29\\358.80 = 340.06\end{array}\] it should be \[ \approx \] I just tried if q=+ and it worked pretty much... (the reason why numbers doesn't fit is because I didn't use more than 2 decimals) I guess you have to try it...

OpenStudy (anonymous):

Is there a button on the scientific calculator I can use? to figure out the sq root thing?

OpenStudy (anonymous):

well I don't know what kind of calculator you have: (In pc) I have\use Microsoft mathematics.... and there I just type q^8= something... and it finds that...

OpenStudy (anonymous):

A scientific one

OpenStudy (amistre64):

your asking how to turn a undo a ^8 right?

OpenStudy (anonymous):

I know I never said that I found a_19 I just used a_9 to see if q=+1.49 works... if it was - there would be 358 =-340 and that's not true... but now to find a19 just substitute into the formula \[\LARGE a_{19}=a_1\cdot r^{19-1} \]

OpenStudy (amistre64):